# Stand-In: A Lightweight and Plug-and-Play Identity Control for Video Generation

### v2

@@ Abstract
Generating high-fidelity human videos that match user-specified identities is important yet challenging in the field of generative AI. Existing methods often rely on an excessive number of training parameters and lack compatibility with other AIGC tools. In this paper, we propose Stand-In, a lightweight and plug-and-play framework for identity preservation in video generation. Specifically, we introduce a conditional image branch into the pre-trained video generation model. Identity control is achieved through restricted self-attentions with conditional position mapping, and can be learned quickly with only 2000 pairs. Despite incorporating and training just $\sim$1% additional parameters, our framework achieves excellent results in video quality and identity preservation, outperforming other full-parameter training methods. Moreover, our framework can be seamlessly integrated for other tasks, such as subject-driven video generation, pose-referenced video generation, stylization, and face swapping.
@@
00footnotetext: * Equal contribution.
Introduction
@@
With the rapid advancement of diffusion models (Ho, Jain, and Abbeel 2020; Podell et al. 2024; Peebles and Xie 2023), video generation (Zheng et al. 2024; Peng et al. 2025; Kong et al. 2024; Hong et al. 2023) has become a pivotal aspect of generative AI. Among its diverse applications, identity-preserving video generation holds profound significance. The goal of this task is to generate high-quality videos that consistently maintain the identity of a given reference image containing a face. It has widespread utility across film, advertising and gaming industries, etc.
Existing methods can be classified into two categories. Traditional methods (He et al. 2024; Yuan et al. 2025b) use an explicit face encoder for identity feature extraction. Recent methods (Hu et al. 2025; Liu et al. 2025) fully train the diffusion transformer. However, face-encoder-based methods lack flexibility and struggle to capture fine facial details essential for high-quality video generation. The full-parameter training methods require a large number of parameters and lack compatibility with other applications. Achieving robust identity preservation in a lightweight and flexible way remains critical yet challenging.
To overcome these limitations, we leverage the pre-trained VAE from the video generation model itself, enabling the conditional image to be mapped directly into the same latent space as the video. This approach naturally utilizes the model’s inherent capabilities to extract rich and detailed facial features, offering a more integrated and effective solution. Furthermore, to achieve a lightweight design and avoid introducing an excessive number of parameters, we employ restricted self-attention with conditional position mapping. On the one hand, it can effectively merge the features of the reference image into the video; on the other hand, it does not alter the architecture of the main video generation model and thus can be used in a plug-and-play manner for other applications, as shown in Figure 1. Experiments demonstrate that our method achieves the highest facial similarity and naturalness in identity-preserving video generation, with the least number of need-to-train parameters, as shown in Figure 2. Thanks to its plug-and-play ability, our framework can be extended to various tasks, including subject-driven generation, video stylization, and face swapping, all while preserving identity consistency. Additionally, by integrating compatibility with VACE (Jiang et al. 2025), our approach significantly enhances facial similarity in pose-guided video generation. Our main contributions can be summarized as follows:
  • To inject identity information without explicit face feature extractors, we introduce a conditional image branch to the video generation model. The image and video branches share information through restricted self-attention with conditional position mapping. With this lightweight design, identity preservation can be learned well only with a small dataset.
Our framework, dataset, and more can be found in our website: https://stand-in-video.github.io/.
Related Work
Video Generation Models
Current video generation models are predominantly built on diffusion frameworks (Ho, Jain, and Abbeel 2020), with a notable evolution in architecture from U-Net-based designs (Blattmann et al. 2023) to DiT-based approaches (Kong et al. 2024; Team 2025; Ma et al. 2025). In the era of U-Net-based diffusion models, text-to-image (T2I) frameworks (Rombach et al. 2022; Podell et al. 2024) were extended to video generation by introducing 3D convolutions and temporal attention  (Blattmann et al. 2023). AnimateDiff (Guo et al. 2024) further advanced this direction by reusing pre-trained text-to-image model weights to leverage their strong spatial generation capabilities by adding temporal layers. Latte (Ma et al. 2025) introduced a spatial-temporal separation mechanism, assigning distinct DiT blocks to process spatial and temporal information independently. This approach was later replaced by 3D full attention mechanisms, which offered more integrated processing. CogVideoX (Yang et al. 2025) and HunyuanVideo (Kong et al. 2024) combined 3D-VAE  (Yu et al. 2024) with MM-DiT (Esser et al. 2024) to enhance video generation capabilities. WAN2.1 (Team 2025) employs a 3D-VAE and adopts a DiT backbone for denoising, injecting semantic prompt information into the diffusion process through cross-attention.
@@
Conventional methods commonly rely on explicit face encoders for facial feature extraction to generate videos with identity-preserving. ID-animator (He et al. 2024) leverages a pre-trained text-to-video diffusion model in conjunction with a lightweight face adapter to encode ID-relevant embeddings from adaptable facial latent queries. ConsistID (Yuan et al. 2025b) aims to maintain identity consistency through frequency decomposition in diffusion transformer. Phantom (Liu et al. 2025) can also preserve identity consistency in the human domain as a unified subject-consistent video generation framework. HunyuanCustom (Hu et al. 2025) is a multi-modal customized video generation framework that emphasizes identity consistency while supporting diverse input modalities. By introducing advanced condition injection mechanisms and identity-preserving strategies, it achieves excellent performance in high-quality video generation. They employed a full fine-tuning for the diffusion transformer, resulting in a large number of trainable parameters.
Method
@@
In this section, we first introduce the overall framework of the proposed method. Next, we detail the restricted self-attention mechanism with conditional position mapping. Finally, we present the data collection process.
Conditional Image Branch
@@
To extract facial features, traditional methods rely on explicit face encoders, which lack flexibility and often fail to preserve fine facial details critical for high-quality reconstruction. In contrast, we propose using the pre-trained VAE from the video generation model.
This strategy maps the conditional image directly into the same latent space as the video and allows us to naturally take advantage of the built-in ability of the pre-trained video generation model to extract rich facial features.
@@
The overall framework is illustrated in Figure 3. We use Wan2.1 14B T2V (Team 2025) as the video generation base model, which adopts a Diffusion Transformer (DiT) architecture. Given a reference image containing a face, we first encode it into the latent space using the pre-trained VAE encoder. The image latents undergo the same patchification and encoding procedures as the video latents.
Then, the image tokens are concatenated with video tokens along the sequence dimension and processed jointly through successive blocks. Finally, image tokens are discarded at the final layer.
Let $s$ denote the denoising timestep in diffusion. To preserve the static nature of the reference image, which serves as a conditioning input rather than undergoing denoising, we maintain its temporal invariance. This is done by fixing its timestep to zero, i.e.,  $s_{ref}=0$.
Now that we have encoded the conditional image into the same feature space as the video, the next challenge is: How can the video features effectively refer to the image information in a way that is lightweight and easy to learn?
Restricted Self-Attention
@@
Figure 4: Design of our restricted self-attention: For the input video and image tokens, we compute their Query, Key, and Value matrices independently. Next, we apply 3D RoPE to the Query and Key matrices. Finally, the image matrices operate independently, while the video Query performs attention using the concatenation of the image and video Key and Value matrices.
In the aforementioned DiT blocks, reference image and video tokens are processed independently through most modules (including layer normalization, cross-attention, and feed-forward networks), except for the self-attention layer. The self-attention layer enables information exchange among all tokens, naturally allowing video tokens to refer to identity information. However, since the reference image serves as a static condition, it should remain unaffected by the dynamic contents of the video. Therefore, to incorporate identity information while preserving its independence, we propose replacing the self-attention layers in DiT with a restricted version that explicitly prevents image queries from attending to video keys.
As shown in Figure 4, for a self-attention layer, we first independently compute the Query, Key, and Value for image and video tokens, denoted as $Q_{I}$, $K_{I}$, $V_{I}$ and $Q_{V}$, $K_{V}$, $V_{V}$ respectively. Then, we concatenate $K_{V}$ with $K_{I}$ and $V_{V}$ with $V_{I}$ for $Q_{V}$. To enhance the model’s ability to utilize identity-related information while preserving its inherent generative robustness, we incorporate Low-Rank Adaptation (LoRA) into the QKV projection of image tokens.
Conditional Position Mapping
To effectively differentiate image and video tokens in the restricted self-attention, we use a specialized conditional position mapping strategy. Specifically, we employ 3D Rotary Positional Embedding (RoPE) (Su et al. 2024), where all tokens associated with the reference image are assigned a distinct and dedicated coordinate space. This ensures clear separation and facilitates precise modeling of interactions between the reference image and video tokens.
@@
This non-overlapping spatial allocation achieves two main goals through geometric separation. By separating the reference tokens from the video coordinate grid, the design naturally reduces false spatial correlations that could lead the model to rely too much on superficial pixel-level matches. At the same time, this separate coordinate system maintains the reference image’s semantic meaning by keeping it as a global identity prior. Consequently, the model is guided to focus on extracting overall semantic features from the reference tokens, rather than treating them as spatially localized patterns that need to align positionally with the video content.
@@
KV Caching
Given that the timestep for the conditional image is fixed at $s_{ref}=0$, its Key and Value matrices remain constant throughout the diffusion denoising process. Therefore, during inference, we can cache $K_{I}$ and $V_{I}$ to accelerate computation. These matrices are computed and stored during the first denoising step, eliminating the need for redundant recalculations in subsequent steps.
Dataset Collection and Processing
Figure 5: Examples from our human-centric video dataset.
We construct a human-centric video dataset containing 2,000 high-resolution sequences from publicly available sources. The dataset guarantees a diverse and comprehensive representation, comprising various ethnic groups, age ranges, gender identities, and a wide array of actions. Using the VILA (Lin et al. 2024) multimodal captioning framework, we automatically generate dense textual annotations for each video, establishing strong text-video alignment.
@@
    Face regions are detected and cropped using RetinaFace(Deng et al. 2020).
@@
    BiSeNet (Yu et al. 2018) is used for face parsing, and the background is replaced with a solid white color to prevent any leakage of background information.
@@
Method Open-Source Face Similarity $\uparrow$ Naturalness $\uparrow$ Prompt Following $\uparrow$
Kling $\times$ 0.410 3.900 19.921
Hailuo $\times$ 0.577 3.750 20.649
Pika-2.1 $\times$ 0.323 3.644 20.649
Vidu-2.0 $\times$ 0.361 3.600 18.998
ID-Animator $\checkmark$ 0.316 3.211 16.677
SkyReels-A2-P14B $\checkmark$ 0.546 3.411 19.110
EchoVideo $\checkmark$ 0.487 3.456 19.263
ConcatID-CogVideoX $\checkmark$ 0.439 3.372 19.359
ConcatID-WAN-AdaLN $\checkmark$ 0.501 3.650 19.671
Hunyuan-Custom $\checkmark$ 0.622 3.367 19.853
ConsistID $\checkmark$ 0.432 3.233 20.552
VACE-P1.3B $\checkmark$ 0.180 3.567 20.591
VACE-1.3B $\checkmark$ 0.223 3.611 20.527
VACE-14B $\checkmark$ 0.647 3.728 19.520
Phantom-1.3B $\checkmark$ 0.440 3.567 20.364
Phantom-14B $\checkmark$ 0.519 3.828 20.476
Stand-In (Ours) $\checkmark$ 0.724 3.922 20.594
@@
Hunyuan-Custom 3.34 2.92
VACE-14B 3.00 3.07
Phantom-14B 2.37 2.92
ConsistID 2.25 2.46
@@
Experiments
Implementation Details
Quantitative Analysis
To evaluate identity preservation and visual quality, we use the two most important and heavily weighted evaluation metrics from the OpenS2V benchmark (Yuan et al. 2025a): facial similarity and naturalness. To evaluate the relevance between generated video and textual description, we use X-CLIP (Ma et al. 2022), a pre-trained video-text multimodal model. Results are shown in Table 1 and Figure 6.

### v3

@@ Abstract
Generating high-fidelity human videos that match user-specified identities is important yet challenging in the field of generative AI. Existing methods often rely on an excessive number of training parameters and lack compatibility with other AIGC tools. In this paper, we propose Stand-In, a lightweight and plug-and-play framework for identity preservation in video generation. Specifically, we introduce a conditional image branch into the pre-trained video generation model. Identity control is achieved through restricted self-attentions with conditional position mapping. Thanks to these designs, which greatly preserve the pretrained prior of the video generation model, our approach is able to outperform other full-parameter training methods in video quality and identity preservation, even with just $\sim$1% additional parameters and only 2000 training pairs. Moreover, our framework can be seamlessly integrated for other tasks, such as subject-driven video generation, pose-referenced video generation, stylization, and face swapping.
@@
footnotetext: Equal Contribution.footnotetext: # Corresponding Author.
1 Introduction
With the rapid advancement of diffusion models [NEURIPS2020_4c5bcfec, podell2023sdxl, Peebles_2023_ICCV], video generation [opensora, opensora2, kong2024hunyuanvideo, hong2022cogvideo] has become a pivotal aspect of generative AI. Among its diverse applications, identity-preserving video generation holds profound significance. The goal of this task is to generate high-quality videos that consistently maintain the identity of a given reference image containing a face. It has widespread utility across film, advertising and gaming industries, etc.
Existing methods can be roughly classified into two categories. Early methods [ID-Animator, consistid] use an explicit face encoder for identity feature extraction, while recent methods [hu2025hunyuancustom, liu2025phantom] fully train the diffusion transformer. However, face-encoder-based methods lack flexibility and struggle to capture fine facial details essential for high-quality video generation. The full-parameter training methods require huge training resources and lack compatibility with other applications. Achieving robust identity preservation in a lightweight way remains critical yet challenging.
To overcome these limitations, we leverage the pre-trained VAE from the video generation model itself, enabling the conditional image to be mapped directly into the same latent space as the video. This approach naturally utilizes the model’s inherent capabilities to extract rich and detailed facial features, offering a more integrated and effective solution. Specifically, we employ restricted self-attention with conditional position mapping to merge the features of the reference image into the video. On the one hand, by preserving the core functionality of self-attention and the pretrained prior of the video generation model, our method achieves the highest facial similarity and naturalness in identity-preserving video generation with only $\sim$1% additional parameters and 2000 training pairs, which is shown in Figure 2. On the other hand, our method does not alter the architecture of the main video generation model and thus can be used in a plug-and-play manner for other applications. As shown in Figure 1, our framework can be extended to various tasks, including subject-driven generation, video stylization, and face swapping, all guaranteeing identity consistency. Additionally, benefit from integrating compatibility with VACE [vace], our approach significantly enhances facial similarity in pose-guided video generation. Our main contributions can be summarized as follows:
  • To inject identity information without explicit face feature extractors, we introduce a conditional image branch to the video generation model. The image and video branches share information through restricted self-attention with conditional position mapping. With these designs, identity preservation can be learned well with a small dataset.
@@ 1 Introduction
2 Related Work
Video Generation
Current video generation models are predominantly built on diffusion frameworks [NEURIPS2020_4c5bcfec], with an evolution in architecture from U-Net-based designs [blattmann2023stablevideodiffusionscaling] to DiT-based approaches [kong2024hunyuanvideo, wan2.1, ma2025latte, DiT, wang2023laviehighqualityvideogeneration, VideoCrafter2, T2V-Turbo]. In the era of U-Net-based diffusion models, text-to-image (T2I) frameworks [rombach2021highresolution, podell2023sdxl] were extended to video generation by introducing 3D convolutions and temporal attention  [blattmann2023stablevideodiffusionscaling]. AnimateDiff [guo2023animatediff] further advanced this direction by reusing pre-trained text-to-image model weights to leverage their strong spatial generation capabilities by adding temporal layers. Latte [ma2025latte] introduced a spatial-temporal separation mechanism, assigning distinct DiT blocks to process spatial and temporal information independently. This approach was later replaced by 3D full attention mechanisms, which offered more integrated processing. CogVideoX [yang2024cogvideox] and HunyuanVideo [kong2024hunyuanvideo] combined 3D-VAE  [yu2023language] with MM-DiT [esser2024scaling] to enhance video generation capabilities. WAN2.1 [wan2.1] employs a 3D-VAE and adopts a DiT backbone for denoising, injecting semantic information into the diffusion process through cross-attention.
@@ 2 Related Work
Prior to the advent of zero-shot identity-preserving algorithms, generating content with consistent identity typically relied on case-by-case fine-tuning.[Dreambooth, Lora, textual_inversion, multi-concept, dreamvideo, customvideo, motionbooth], In contrast, training-free identity-preserving image generation methods[ye2023ip, shi2024instantbooth, guo2025pulid, wang2024instantid, he2024uniportrait, sara2025ipcompose, qian2024omni, han2024face, mou2025dreamounifiedframeworkimage] offer a zero-shot personalization solution by integrating identity features into pre-trained foundation models. These methods typically introduce parameterized plug-in modules or adapters to adjust and inject identity features into the generation process. A popular solution is to use a Face Encoder to extract face embeddings and inject them into the generation process via cross-attention. For example, InstantID[wang2024instantid] and PuLID[guo2025pulid] can generate high-quality, identity-preserving images.
@@
In the field of identity-preserving video generation. Early methods commonly rely on explicit face encoders for facial feature extraction to generate videos with identity-preserving. ID-animator [ID-Animator] leverages a pre-trained text-to-video diffusion model in conjunction with a lightweight face adapter to encode ID-relevant embeddings from adaptable facial latent queries. ConsistID [consistid] aims to maintain identity consistency through frequency decomposition in diffusion transformer. Phantom [liu2025phantom] can also preserve identity consistency in the human domain as a unified subject-consistent video generation framework. HunyuanCustom [hu2025hunyuancustom] is a multi-modal customized video generation framework that emphasizes identity consistency while supporting diverse input modalities. By introducing advanced condition injection mechanisms and identity-preserving strategies, it achieves excellent performance in high-quality video generation. They employed a full fine-tuning for the diffusion transformer, requiring huge training resources.
3 Method
@@
In this section, we first introduce the overall framework of the proposed method in Section 3.1. Next, we detail the restricted self-attention mechanism in Section 3.2 and conditional position mapping in Section 3.3. Finally, we present the data collection process in Section 3.4.
3.1 Overall Framework
@@
To extract facial features, early methods [ID-Animator, consistid] rely on explicit face encoders, which lack flexibility and often fail to preserve fine facial details critical for high-quality reconstruction. In contrast, we propose using the pre-trained VAE from the video generation model. This strategy maps the conditional image directly into the same latent space as the video and allows us to naturally take advantage of the built-in ability of the pre-trained video generation model to extract rich facial features.
@@
The overall framework is illustrated in Figure 3. We use Wan2.1-14B-T2V [wan2.1] as the video generation base model, which adopts a Diffusion Transformer (DiT) architecture. Given a reference image containing a face, we first encode it into the latent space using the pre-trained VAE encoder. The image latents undergo the same patchification and encoding procedures as the video latents. Then, the image tokens are concatenated with video tokens along the sequence dimension and processed jointly through successive blocks. Finally, image tokens are discarded at the final layer.
[…]
3.2 Restricted Self-Attention
@@
Figure 4: Design of our Restricted Self-Attention: For the input video and image tokens, we compute their Query, Key, and Value matrices independently. Next, we apply Conditional Position Mapping to the Query and Key matrices. Finally, the image matrices operate independently, while the video Query performs attention using the concatenation of the image and video Key and Value matrices.
In the aforementioned DiT blocks, reference image and video tokens are processed independently through most modules (including layer normalization, cross-attention, and feed-forward networks), except for the self-attention layer. The self-attention layer enables information exchange among all tokens, naturally allowing video tokens to refer to identity information. One intuitive solution is to direct concatenate the reference-image tokens with the generated video tokens, and pass them through the Vanilla Self-Attention. However, this approach has two main drawbacks. First, since the reference image serves as a static condition, it should remain unaffected by the dynamic contents of the video. Vanilla Self-Attention lets image queries attend to video contents, making it challenging to maintain the identity. Second, this joint self-attention provides no guarantee that video tokens will actually refer to the image tokens. The model may ignore the reference image and generate scenes without the target identity, which is shown in Figure 5.
Therefore, to incorporate identity information while preserving its independence, we propose replacing the vanilla self-attention layers in DiT with a restricted version that explicitly prevents image queries from attending to video keys. As shown in Figure 4, for a self-attention layer, we first independently compute the Query, Key, and Value for image and video tokens, denoted as $Q_{I}$, $K_{I}$, $V_{I}$ and $Q_{V}$, $K_{V}$, $V_{V}$ respectively. Then, we concatenate $K_{V}$ with $K_{I}$ and $V_{V}$ with $V_{I}$ for $Q_{V}$. To enhance the model’s ability to utilize identity-related information while preserving its inherent generative robustness, we incorporate Low-Rank Adaptation (LoRA) into the QKV projection of image tokens. For analysis, we visualize attention maps specifically for video queries attending to the reference image by averaging attention maps across all DiT blocks. As shown in Figure 4, in contrast to the vanilla baseline, our Restricted Self-Attention concentrates attention on facial regions of the reference image and yields prompt-faithful frames that preserve the subject’s identity.
Given that the timestep for the conditional image is fixed at $s_{ref}=0$, its Key and Value matrices remain constant throughout the diffusion denoising process. Therefore, during inference, we can cache $K_{I}$ and $V_{I}$ to accelerate computation, named KV Caching. These matrices are computed and stored during the first denoising step, eliminating the need for redundant recalculations in subsequent steps.
3.3 Conditional Position Mapping
To effectively differentiate image and video tokens in the restricted self-attention, we use a specialized conditional position mapping strategy. Specifically, we employ 3D Rotary Positional Embedding (RoPE) [RoPE], where all tokens associated with the reference image are assigned a distinct and dedicated coordinate space. This ensures clear separation and facilitates precise modeling of interactions between the reference image and video tokens.
@@ 3.3 Conditional Position Mapping
Figure 5: Effect of Restricted Self-Attention (RSA). Given the reference image and the prompt (left column), we visualize attention map to reference-image tokens. Under Vanilla Self-Attention (top row), attention diffuses into background regions and the output skews toward a garden scene. With RSA (bottom row), attention concentrates on facial regions, maintaining the subject’s identity.
@@
Figure 6: Effect of Conditional Position Mapping (CPM). Compared with Shared Postion Mapping, our CPM, where the reference tokens are mapped to a disjoint spatial space, better preserves the pretrained positional prior and yields more stable scenes.
This non-overlapping spatial allocation achieves two main goals through geometric separation. By separating the reference tokens from the video coordinate grid, the design naturally reduces false spatial correlations and better preserves the backbone’s pretrained positional prior. Compared with Shared Position Mapping, our Conditional Position Mapping generates more reliable video with more stable scene layouts, which is shown in Figure 6. At the same time, this separate coordinate system maintains the reference image’s semantic meaning by keeping it as a global identity prior. Consequently, the model is guided to focus on extracting overall semantic features from the reference tokens, rather than treating them as spatially localized patterns that need to align positionally with the video content.
3.4 Dataset Collection and Processing
We construct a human-centric video dataset containing 2,000 high-resolution sequences from publicly available sources. The dataset guarantees a diverse and comprehensive representation, comprising various ethnic groups, age ranges, gender identities, and a wide array of actions. Using the VILA [lin2024vila] multimodal captioning framework, we automatically generate dense textual annotations for each video, establishing strong text-video alignment.
@@ 3.4 Dataset Collection and Processing
    Faces are detected and cropped using RetinaFace[Deng2020CVPR].
@@
    BiSeNet [BiseNet] is used for face parsing, and the background is replaced with a solid white color to prevent any leakage of background information.
Figure 7: Examples from our human-centric video dataset.
@@
Method
Face-
Similarity
$\uparrow$
Naturalness $\uparrow$
Prompt-
Following
$\uparrow$
Closed-Source Methods
Kling 0.410 3.900 19.921
Hailuo 0.577 3.750 20.649
Pika-2.1 0.323 3.644 20.649
Vidu-2.0 0.361 3.600 18.998
Open-Source Methods
ID-Animator [ID-Animator] 0.316 3.211 16.677
SkyReels-A2-P14B [SkyReels] 0.546 3.411 19.110
EchoVideo [EchoVideo] 0.487 3.456 19.263
ConcatID-CogVideoX [Concat-ID] 0.439 3.372 19.359
ConcatID-WAN [Concat-ID] 0.501 3.650 19.671
Hunyuan-Custom [HunyuanCustom] 0.622 3.367 19.853
ConsistID [consistid] 0.432 3.233 20.552
VACE-P1.3B [vace] 0.180 3.567 20.591
VACE-1.3B [vace] 0.223 3.611 20.527
VACE-14B [vace] 0.647 3.728 19.520
Phantom-1.3B [liu2025phantom] 0.440 3.567 20.364
Phantom-14B [liu2025phantom] 0.519 3.828 20.476
Stand-In (Ours) 0.724 3.922 20.594
4 Experiments
4.1 Implementation Details
@@
Hunyuan-Custom[HunyuanCustom] 3.34 2.92
VACE-14B [vace] 3.00 3.07
Phantom-14B [liu2025phantom] 2.37 2.92
ConsistID [consistid] 2.25 2.46
@@
Figure 9: Our model generalizes to unseen ordinary individuals across diverse ethnicities and age groups, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs.
@@
4.2 Quantitative Analysis
To evaluate identity preservation and visual quality, we use the two most important and heavily weighted evaluation metrics from the OpenS2V benchmark [yuan2025opens2vnexusdetailedbenchmarkmillionscale]: facial similarity and naturalness. To evaluate the relevance between generated video and textual description, we use X-CLIP [xclip], a pre-trained video-text multimodal model. Results are shown in Table 1 and Figure 8. As shown in Figure 9, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs, our model also demonstrates strong generalization across individuals of diverse ethnicities and age groups.